OpenStudy (anonymous):

Help?... Answer the following One angle in a rhombus measures 65 degrees which of the following is the measure of another angle in the rhombus? 25 Degrees 55 Degrees 85 Degrees 115 Degrees

OpenStudy (anonymous):

85, maybe?

OpenStudy (anonymous):

Can you like explain?

OpenStudy (anonymous):

Probably not. I'm bad at math. I'm sorry.

OpenStudy (anonymous):

Oh its okay..

OpenStudy (anonymous):

115

OpenStudy (anonymous):

How did you get 115?

OpenStudy (anonymous):

since in a rhombus the opposite angles are congruent and it is a quadrilateral so the interior angles need to add up to 360. 65*2=130 360-130=230 230/2=115

OpenStudy (anonymous):

Oh okay thanks.

OpenStudy (anonymous):

can u help me out on a q?

OpenStudy (anonymous):

I can try.

OpenStudy (anonymous):

i tagged u